Prove that the union of two closed sets is a closed set. Give an example to show that

union of an infinite number of closed sets need not be a closed set.


1
Expert's answer
2021-05-31T15:59:23-0400

Solution:

Let "(X, \mathcal{T})" be a topological space.

By definition, for any subset "A \subset X" we have

A is open "\Longleftrightarrow A \in \mathcal{T}" and A is closed "\Longleftrightarrow C(A)" is open.

a)"\emptyset" is closed: Because "\emptyset=C(X)" and "X \in \mathcal{T}"

b) X is closed: Because "X=C(\emptyset)" and "\emptyset \in \mathcal{T}"

c) A finite union of closed sets is closed: Let "A_{1}, \ldots, A_{N}" be closed subsets in X. Then "C\left(A_{1}\right), \ldots, C\left(A_{N}\right)" are open by definition. Furthermore, we have

"C\left(\bigcup_{n=1}^{N} A_{n}\right)=\bigcap_{n=1}^{N} C\left(A_{n}\right)"

from set theory, and this is open. So its complement

"\bigcup_{n=1}^{N} A_{n}" is closed by definition, as desired.

d) An arbitrary intersection of closed sets is closed: Let "\left\{A_{\alpha}\right\}_{\alpha \in I}" be a collection of closed subsets in X. Then "C\left(A_{\alpha}\right)" is open for every \alpha \in I by definition. Furthermore, we have

"C\left(\bigcap_{\alpha \in I} A_{\alpha}\right)=\bigcup_{\alpha \in I} C\left(A_{\alpha}\right)"

from set theory, and this is open. So its complement

"\bigcap_{\alpha \in I} A_{a}"

is closed by definition, as desired.

From all these parts, we showed that the union of two closed sets is a closed set.


Example:

Define "A_{n}=\left[\frac{1}{n}, 1-\frac{1}{n}\right]" for n>1.

Then, obviously each "A_{n}" is closed, but "\bigcup A_{n}=(0,1)"

which is open and not closed.